# Effective Embodied Representation

### v2

@@ 1 Introduction
Vision systems have made remarkable progress in understanding 2D images (He et al., 2020; Chen et al., 2020a; He et al., 2022; Feichtenhofer et al., 2022; Tong et al., 2022; Yang et al., 2023; Oquab et al., 2023; Radford et al., 2021; Fang et al., 2023b; Chen et al., 2024). However, achieving true visual intelligence necessitates a comprehensive understanding of the 3D world. This is crucial for embodied AI, where agents must perceive, reason, and interact with complex 3D environments.
@@
To thoroughly validate our assumption and method, we collect 268 embodied tasks across 8 simulators using various policy methods. To our knowledge, this represents the largest scale of embodied evaluation to date. Previous work, such as R3M (Nair et al., 2022) and VC-1 (Majumdar et al., 2023), evaluated fewer than 20 tasks, potentially leading to incomplete or biased conclusions. Our evaluation spans both single-task and language-conditioned multi-task learning. We compare over 10 state-of-the-art representation learning methods, categorized as embodied-specific (Nair et al., 2022; Majumdar et al., 2023; Radosavovic et al., 2023), vision-centric (Oquab et al., 2023; Chen et al., 2021; He et al., 2022), and multi-modal (Radford et al., 2021; Fang et al., 2023b; Chen et al., 2024). Our method consistently outperforms others, underscoring the importance of 3D spatial awareness for embodied AI. Notably, multi-modal models like CLIP (Radford et al., 2021), consistently perform poorly. This holds even the vision-language model scales the ViT to 6B parameters (Chen et al., 2024). Through a camera pose estimation task and feature map visualization, we demonstrate that SPA has learned superior 3D spatial understanding. Further, we find that 3D awareness shows a positive correlation with embodied performance. Finally, we conduct several real-world tasks, where SPA also demonstrates superior performance. Our contribution can be summarized as follows.
@@ 4 Training and Implementation Details
In this section, we present the step-by-step implementation and training of our SPA model. We first pre-train a ViT-base (ViT-B) backbone using a small dataset and evaluate this model on the VC-1 benchmark to examine the effects of hyperparameters (Sec. 1). We then compile several multi-view datasets, training ViT-B models on each to assess the impact of different datasets (Sec. 4.2). Finally, we integrate all factors and scale up both data and model size to train the strongest version of SPA using a ViT-large (ViT-L) backbone (Sec. 4.3). More details can be found in Appendix C.
4.1 Hyperparameter Investigation
Table 1: Influence of mask ratio and loss components. C., D., and S. denote the color, depth, and semantic loss components, respectively.
Mask Ratio Loss VC-1 Benchmark Mean S.R.
C. D. S. AD MW DMC TF
0.00 53.3$\pm$4.6 88.5$\pm$5.7 57.5$\pm$2.6 74.1$\pm$0.6 70.36
0.25 52.7$\pm$3.1 89.6$\pm$4.5 57.6$\pm$3.0 70.4$\pm$1.7 70.17
0.50 53.3$\pm$4.2 88.8$\pm$1.6 60.1$\pm$3.1 72.6$\pm$0.7 71.18
0.75 51.3$\pm$1.2 88.0$\pm$3.5 61.1$\pm$3.5 73.0$\pm$0.8 71.01
0.95 51.3$\pm$1.2 85.6$\pm$4.0 62.5$\pm$5.3 73.1$\pm$0.2 70.67
0.50 51.3$\pm$1.2 90.9$\pm$3.3 58.8$\pm$5.6 71.5$\pm$1.0 71.01
0.50 52.0$\pm$2.0 89.3$\pm$3.3 53.9$\pm$4.3 70.9$\pm$1.3 68.71
0.50 52.7$\pm$3.1 88.0$\pm$4.5 61.5$\pm$3.4 71.6$\pm$1.2 71.16
@@ 5.1 Overall Comparisons (Q1, Q2)
Method Vision-Centric Multi-Modal Embodied-Specific
MoCoV3 (Chen et al., 2020b) MAE (He et al., 2022) DINOV2 (Oquab et al., 2023) CLIP (Radford et al., 2021) EVA (Fang et al., 2023b) InternViT-300M (Chen et al., 2024) InternViT-6B (Chen et al., 2024) MVP (Radosavovic et al., 2023) VC-1 (Majumdar et al., 2023) SPA (Ours)
@@
Input Size 224 224 224 224 224 448 224 256 224 224
Patch Size 16 16 14 14 14 14 14 16 16 16
#Param. 303M 303M 303M 303M 303M 303M 5.9B 303M 303M 303M
#Frames 1.28M 1.28M 1.2B 400M 14M 5.0B 5.0B 4.5M 5.6M 3.8M
Table 4: Comparison of different representation learning methods. ‘OOM’ indicates InternViT-6B encountered an out-of-memory error during evaluation. The best and second-best results are bolded and underlined respectively. The number in parentheses denotes the number of tasks.
@@
Baselines. We evaluate 9 state-of-the-art representation learning models, all using the same ViT-L backbone, categorized into vision-centric, multi-modal, and embodied-specific. This also includes a comparison with a 6B-parameter multi-modal model (Chen et al., 2024). Details of the models are summarized in Tab. 3. The results on each benchmark are shown in Tab. 4. For detailed results on each task and each random seed, please refer to Appendix D. We also have visualized the performance radar chart and the per-task rank distributions in Fig. 1.
@@
Table 5: Additional comparisons of ViT-base models on VC-1 benchmarks.
@@ 5.3 Study on 3D Awareness of SPA (Q3)
Table 6: Zero-shot camera pose estimation. SPA demonstrates strong 3D awareness.
@@ 5.4 Real-World Experiments (Q4)
Table 8: Real-world experiment results.
@@ 6 Related Work
Representation Learning for Computer Vision. Recent advances in computer vision have increasingly focused on unsupervised and self-supervised learning to utilize large amounts of unlabeled data. Techniques like contrastive learning (Chen et al., 2020a; 2021; b; He et al., 2020), masked autoencoders (He et al., 2022; Feichtenhofer et al., 2022; Bachmann et al., 2022; Tong et al., 2022; Wang et al., 2023), and self-distillation (Caron et al., 2021; Oquab et al., 2023; Ranzinger et al., 2024) have shown that effective representations can be learned without supervision. Moreover, multi-modal pre-training approaches (Radford et al., 2021; Fang et al., 2023b; Chen et al., 2024) leverage language to learn more comprehensive representations. These developments have significantly improved transfer learning capabilities while also displaying zero-shot abilities.
Representation Learning for Embodied AI. Inspired by computer vision, recent work has applied methods such as contrastive learning (Nair et al., 2022; Yang et al., 2023) and masked autoencoders (Radosavovic et al., 2023; Majumdar et al., 2023; Karamcheti et al., 2023; Yang et al., 2024b) to embodied AI. However, these approaches often overlook the unique challenges of embodied tasks, focusing primarily on semantic knowledge. However, many of these methods primarily adapt techniques from computer vision and focus predominantly on learning semantic knowledge, often overlooking the unique requirements of embodied AI tasks. Prior studies (Zhu et al., 2024; Ze et al., 2024; Wang et al., 2024b; a) have highlighted the importance of 3D spatial structure for improving robotic learning, though point clouds are difficult to scale. Despite this, point clouds are challenging to obtain and encode, limiting their scalability. In this work, we propose a spatial hypothesis: while semantic understanding is crucial, 3D spatial awareness is even more important, and we demonstrate how a standard 2D backbone can integrate 3D spatial awareness.
@@
7 Conclusion and Limitations
In this work, we propose that 3D spatial awareness is crucial for embodied AI and introduce SPA, a novel framework that pre-trains a standard ViT backbone with 3D spatial awareness. To validate our hypothesis, we conduct the largest-scale embodied evaluation to date, over 15 times larger than previous studies. Our experiments demonstrate the clear superiority of SPA and highlight the importance of 3D awareness. Despite strong results across simulated and real robotic tasks, limitations remain. Our evaluation is currently restricted to imitation learning (specifically behavior cloning), and exploring SPA’s performance in other settings, such as reinforcement learning, presents an exciting future direction. Additionally, SPA currently focuses on static multi-view scenes; extending it to dynamic, temporal scenarios could enhance its generality. Lastly, while we use the ViT encoder for fair comparison, the volume decoder’s multi-view interaction knowledge could be leveraged in policy learning, offering further potential for improvement.
@@ Appendix B Evaluation Setups
Here we detail the setups of our large-scale evaluation in Sec. 3.
@@ B.1 Single-Task Benchmarks
VC-1 (Majumdar et al., 2023). This benchmark includes several simulators. We selected four: Adroit (Kumar, 2016), Meta-World (Yu et al., 2020), DMControl (Tunyasuvunakool et al., 2020), and TriFinger (Wüthrich et al., 2020). The Adroit subset focuses on dexterous manipulation with 2 tasks. The Meta-World subset addresses two-finger gripper manipulation with 5 tasks. The DMControl subset is for locomotion control, also with 5 tasks. The TriFinger subset targets three-finger manipulation with 2 tasks. For all tasks, we use a 3-layer MLP as the policy network for each single-task training, following the original implementation. Each task is trained with 100 demonstrations, except for 25 on Meta-World, and evaluated 50 times using the specific seeds 100, 200, and 300. The [CLS] token of a frozen pre-trained ViT is used as the observation feature. All hyper-parameters are kept the same with the original implementation.
@@
Meta-World (Yu et al., 2020). This benchmark comprises a series of tasks in which an agent directs a Sawyer robot arm to manipulate objects in a tabletop environment. We selected 48 tasks, encompassing easy, medium, and hard levels. We implemented the Diffusion Policy (Chi et al., 2023) on this benchmark and adhered to the setup in Ze et al. (2024) to generate 10 demonstrations for each single-task training, followed by evaluation through 20 rollouts. The average results across three fixed seeds (100, 200, 300) are reported. The [CLS] token from a frozen pre-trained ViT serves as the observation feature.

### v3

@@ 1 Introduction
Vision systems have made remarkable progress in understanding 2D images (He et al., 2020; Chen et al., 2020a; He et al., 2022; Feichtenhofer et al., 2022; Tong et al., 2022; Yang et al., 2023; Oquab et al., 2023; Radford et al., 2021; Fang et al., 2023b; Chen et al., 2024b). However, achieving true visual intelligence necessitates a comprehensive understanding of the 3D world. This is crucial for embodied AI, where agents must perceive, reason, and interact with complex 3D environments.
@@
To thoroughly validate our assumption and method, we collect 268 embodied tasks across 8 simulators using various policy methods. To our knowledge, this represents the largest scale of embodied evaluation to date. Previous work, such as R3M (Nair et al., 2022) and VC-1 (Majumdar et al., 2023), evaluated fewer than 20 tasks, potentially leading to incomplete or biased conclusions. Our evaluation spans both single-task and language-conditioned multi-task learning. We compare over 10 state-of-the-art representation learning methods, categorized as embodied-specific (Nair et al., 2022; Majumdar et al., 2023; Radosavovic et al., 2023), vision-centric (Oquab et al., 2023; Chen et al., 2021; He et al., 2022), and multi-modal (Radford et al., 2021; Fang et al., 2023b; Chen et al., 2024b). Our method consistently outperforms others, underscoring the importance of 3D spatial awareness for embodied AI. Notably, multi-modal models like CLIP (Radford et al., 2021), consistently perform poorly. This holds even the vision-language model scales the ViT to 6B parameters (Chen et al., 2024b). Through a camera pose estimation task and feature map visualization, we demonstrate that SPA has learned superior 3D spatial understanding. Further, we find that 3D awareness shows a positive correlation with embodied performance. Finally, we conduct several real-world tasks, where SPA also demonstrates superior performance. Our contribution can be summarized as follows.
@@ 4 Training and Implementation Details
In this section, we present the implementation and training of our SPA model. We first compile several multi-view datasets, training ViT-B models on each to assess the impact of different datasets (Sec. 4.1). Finally, we integrate all factors and scale up both data and model size to train the strongest version of SPA using a ViT-large (ViT-L) backbone (Sec. 4.2). More details can be found in Appendix C.
@@ 5.1 Overall Comparisons (Q1, Q2)
Method Vision-Centric Multi-Modal Embodied-Specific Distilled AM-RADIO (Ranzinger et al., 2024)
MoCoV3 (Chen et al., 2020b) MAE (He et al., 2022) DINOV2 (Oquab et al., 2023) CLIP (Radford et al., 2021) EVA (Fang et al., 2023b) InternViT-300M (Chen et al., 2024b) InternViT-6B (Chen et al., 2024b) MVP (Radosavovic et al., 2023) VC-1 (Majumdar et al., 2023) SPA (Ours)
@@
Input Size 224 224 224 224 224 448 224 256 224 224 dynamic
Patch Size 16 16 14 14 14 14 14 16 16 16 16
#Param. 303M 303M 303M 303M 303M 303M 5.9B 303M 303M 303M 653M
#Frames 1.28M 1.28M 1.2B 400M 14M 5.0B 5.0B 4.5M 5.6M 3.8M 1.4B
Table 3: Comparison of different representation learning methods. ‘OOM’ indicates an out-of-memory error during evaluation. The best and second-best results are bolded and underlined respectively. The number in parentheses denotes the number of tasks. S.R. denotes ‘Success Rate’.
@@
Baselines. We evaluate 9 SOTA representation learning models, all using ViT-L backbone, categorized into vision-centric, multi-modal, and embodied-specific. This also includes a 6B multi-modal model (Chen et al., 2024b). The vision-centric methods are originally from the vision community; the multi-modal methods are typically CLIP-style language-image pre-trained models and are used specifically for VLMs; the embodied-specific methods are designed and pre-trained specifically for embodied AI tasks. Details are summarized in Tab. 2. The results on each benchmark are shown in Tab. 3. For detailed results on each task and each random seed, please refer to Appendix D. We also have visualized the performance radar chart and the per-task rank distributions in Fig. 1.
@@
Table 4: Additional comparisons of ViT-base models. S.R. denotes ‘Success Rate’.
@@ 5.3 Study on 3D Awareness of SPA (Q3)
Table 5: Zero-shot camera pose estimation. Trans. and Rot. denote ‘translation’ and ‘rotation’ errors respectively. The detailed metrics on the error calculation are listed in Appendix E.
@@
5.4 Hyperparameter Investigation
Table 7: Mask ratio and loss components. C., D., S. denote color, depth, and semantic.
Mask Ratio Loss VC-1 Benchmark Mean S.R.
C. D. S. AD MW DMC TF
0.00 53.3$\pm$4.6 88.5$\pm$5.7 57.5$\pm$2.6 74.1$\pm$0.6 70.36
0.25 52.7$\pm$3.1 89.6$\pm$4.5 57.6$\pm$3.0 70.4$\pm$1.7 70.17
0.50 53.3$\pm$4.2 88.8$\pm$1.6 60.1$\pm$3.1 72.6$\pm$0.7 71.18
0.75 51.3$\pm$1.2 88.0$\pm$3.5 61.1$\pm$3.5 73.0$\pm$0.8 71.01
0.95 51.3$\pm$1.2 85.6$\pm$4.0 62.5$\pm$5.3 73.1$\pm$0.2 70.67
0.50 51.3$\pm$1.2 90.9$\pm$3.3 58.8$\pm$5.6 71.5$\pm$1.0 71.01
0.50 52.0$\pm$2.0 89.3$\pm$3.3 53.9$\pm$4.3 70.9$\pm$1.3 68.71
0.50 52.7$\pm$3.1 88.0$\pm$4.5 61.5$\pm$3.4 71.6$\pm$1.2 71.16
@@ 5.5 Real-World Experiments (Q4)
Table 8: Real-world experiment results. S.R. denotes ‘Success Rate’.
@@ 6 Related Work
Representation Learning for Computer Vision. Recent advances in computer vision have increasingly focused on unsupervised and self-supervised learning to utilize large amounts of unlabeled data. Techniques like contrastive learning (Chen et al., 2020a; 2021; b; He et al., 2020), masked autoencoders (He et al., 2022; Feichtenhofer et al., 2022; Bachmann et al., 2022; Tong et al., 2022; Wang et al., 2023), and self-distillation (Caron et al., 2021; Oquab et al., 2023; Ranzinger et al., 2024) have shown that effective representations can be learned without supervision. Moreover, multi-modal pre-training approaches (Radford et al., 2021; Fang et al., 2023b; Chen et al., 2024b) leverage language to learn more comprehensive representations. These developments have significantly improved transfer learning capabilities while also displaying zero-shot abilities.
Representation Learning for Embodied AI. Recent advances in embodied AI representation learning, inspired by computer vision, have applied techniques such as contrastive (Nair et al., 2022; Yang et al., 2023) and masked autoencoders (Radosavovic et al., 2023; Majumdar et al., 2023; Karamcheti et al., 2023; Yang et al., 2024b) to embodied AI. However, these approaches often emphasize semantic learning while overlooking the specific needs of embodied AI tasks. In this work, we propose a spatial hypothesis specifically for embodied AI representation learning, and we demonstrate how a standard 2D backbone can integrate 3D spatial awareness.
3D Robot Learning and 3D-Aware Computer Vision. Prior work in 3D robot learning has often relied on explicit 3D input (Zhu et al., 2024; Ze et al., 2024; Wang et al., 2024b; a; Shridhar et al., 2023; Chen et al., 2023), or lifting 2D features into 3D spaces (Ke et al., 2024; Goyal et al., 2024), providing a strong foundation for our spatial hypothesis. Given the scalability challenges of explicit 3D observations, some computer vision research has explored integrating 3D spatial awareness into 2D backbones (Yang et al., 2024a; Zhu et al., 2023b; Yue et al., 2025; Zhang et al., 2024). To the best of our knowledge, SPA is the first to systematically investigate this approach in embodied AI.
@@
7 Conclusion, Limitations, and Future Work
In this work, we propose that 3D spatial awareness is crucial for embodied AI and introduce SPA, a novel framework that pre-trains a standard ViT backbone with 3D spatial awareness. To validate our hypothesis, we conduct the largest-scale embodied evaluation to date, over 15 times larger than previous studies. Our experiments demonstrate the clear superiority of SPA and highlight the importance of 3D awareness. Despite strong results across simulated and real robotic tasks, limitations remain. Our evaluation is currently restricted to imitation learning (specifically behavior cloning), and exploring SPA’s performance in other settings, such as reinforcement learning, presents an exciting future direction. Incorporating SPA into VLMs to enhance their performance on spatial aware tasks (Chen et al., 2024a; Majumdar et al., 2024) also represents an exciting research area. Additionally, SPA currently focuses on static multi-view scenes; extending it to dynamic, temporal scenarios could enhance its generality. Lastly, while we use the ViT encoder for fair comparison, the volume decoder’s multi-view interaction knowledge could be leveraged in policy learning, offering further potential for improvement.
Acknowledgments
This work is supported by the National Key R&D Program of China (NO.2022ZD0160102) and Shanghai Artificial Intelligence Laboratory.
@@ Appendix B Evaluation Setups
Here we detail the setups of our large-scale evaluation in Sec. 3. For the detailed visualizations of each task, we recommend the readers to read the original simulator’s or benchmark’s dataset.
@@ B.1 Single-Task Benchmarks
VC-1 (Majumdar et al., 2023). This benchmark includes several simulators. We selected four: Adroit (Kumar, 2016), Meta-World (Yu et al., 2020), DMControl (Tunyasuvunakool et al., 2020), and TriFinger (Wüthrich et al., 2020). The Adroit subset focuses on dexterous manipulation with 2 tasks: Relocate and Pen. The Meta-World subset addresses two-finger gripper manipulation with 5 tasks: Button Press Topdown, Drawer Open, Bin Picking, Hammer, and Assembly. The DMControl subset is for locomotion control, also with 5 tasks: Walker Stand, Walker Walk, Reacher Easy, Cheetah Run, and Finger Spin. The TriFinger subset targets three-finger manipulation with 2 tasks: Reach Cube and Move Cube. For all tasks, we use a 3-layer MLP as the policy network for each single-task training, following the original implementation. Each task is trained with 100 demonstrations, except for 25 on Meta-World, and evaluated 50 times using the specific seeds 100, 200, and 300. The [CLS] token of a frozen pre-trained ViT is used as the observation feature. All hyper-parameters are kept the same with the original implementation.
@@
Meta-World (Yu et al., 2020). This benchmark comprises a series of tasks in which an agent directs a Sawyer robot arm to manipulate objects in a tabletop environment. We selected 48 tasks, encompassing easy, medium, and hard levels. We implemented the Diffusion Policy (Chi et al., 2023) on this benchmark and adhered to the setup in Ze et al. (2024) to generate 10 demonstrations for each single-task training, followed by evaluation through 20 rollouts. The average results across three fixed seeds (100, 200, 300) are reported. The [CLS] token from a frozen pre-trained ViT serves as the observation feature. The 48 tasks include: Button Press Wall, Door Close, Door Unlock, Drawer Close, Drawer Open, Faucet Close, Plate Slide, Plate Slide Back, Plate Slide Side, Window Close, Basketball, Bin Picking, Box Close, Coffee Push, Assembly, Disassemble, Push Wall, Shelf Place, Door Open, Button Press, Sweep Into, Door Lock, Reach Wall, Hammer, Stick Push, Button Press Topdown, Handle Press Side, Plate Slide Back Side, Sweep, Button Press Topdown Wall, Handle Press, Push, Coffee Pull, Dial Turn, Reach, Coffee Button, Pick Place Wall, Stick Pull, Hand Insert, Peg Insert Side, Pick Place, Faucet Open, Push Back, Lever Pull, Handle Pull, Soccer, Window Open, and Pick Out Of Hole.
@@ B.2 Language-Conditioned Multi-Task Benchmarks
The 35 tasks in Group 1 include: Basketball In Hoop, Put Rubbish In Bin, Meat Off Grill, Meat On Grill, Slide Block To Target, Reach And Drag, Take Frame Off Hanger, Water Plants, Hang Frame On Hanger, Wipe Desk, Stack Blocks, Reach Target, Push Button, Lamp On, Toilet Seat Down, Close Laptop Lid, Open Box, Open Drawer, Pick Up Cup, Turn Tap, Take Usb Out Of Computer, Play Jenga, Insert Onto Square Peg, Take Umbrella Out Of Umbrella Stand, Insert Usb In Computer, Straighten Rope, Turn Oven On, Change Clock, Close Microwave, Close Fridge, Close Grill, Open Grill, Unplug Charger, Press Switch, and Take Money Out Safe.
The 36 tasks in Group 2 include: Change Channel, Tv On, Push Buttons, Stack Wine, Scoop With Spatula, Place Hanger On Rack, Move Hanger, Sweep To Dustpan, Take Plate Off Colored Dish Rack, Screw Nail, Take Shoes Out Of Box, Slide Cabinet Open And Place Cups, Lamp Off, Pick And Lift, Take Lid Off Saucepan, Close Drawer, Close Box, Phone On Base, Toilet Seat Up, Put Books On Bookshelf, Beat The Buzz, Stack Cups, Put Knife On Chopping Board, Place Shape In Shape Sorter, Take Toilet Roll Off Stand, Put Umbrella In Umbrella Stand, Setup Checkers, Open Window, Open Wine Bottle, Open Microwave, Put Money In Safe, Open Door, Close Door, Open Fridge, Open Oven, and Plug Charger In Power Supply.
@@
The 10 tasks in LIBERO-Spatial are:
1. Pick up the black bowl between the plate and the ramekin and place it on the plate.
2. Pick up the black bowl next to the ramekin and place it on the plate.
3. Pick up the black bowl from table center and place it on the plate.
4. Pick up the black bowl on the cookie box and place it on the plate.
5. Pick up the black bowl in the top drawer of the wooden cabinet and place it on the plate.
6. Pick up the black bowl on the ramekin and place it on the plate.
7. Pick up the black bowl next to the cookie box and place it on the plate.
8. Pick up the black bowl on the stove and place it on the plate.
9. Pick up the black bowl next to the plate and place it on the plate.
10. Pick up the black bowl on the wooden cabinet and place it on the plate.
The 10 tasks in LIBERO-Object are:
1. Pick up the alphabet soup and place it in the basket.
2. Pick up the cream cheese and place it in the basket.
3. Pick up the salad dressing and place it in the basket.
4. Pick up the BBQ sauce and place it in the basket.
5. Pick up the ketchup and place it in the basket.
6. Pick up the tomato sauce and place it in the basket.
7. Pick up the butter and place it in the basket.
8. Pick up the milk and place it in the basket.
9. Pick up the chocolate pudding and place it in the basket.
10. Pick up the orange juice and place it in the basket.
The 10 tasks in LIBERO-Goal are:
1. Open the middle drawer of the cabinet.
2. Put the bowl on the stove.
3. Put the wine bottle on top of the cabinet.
4. Open the top drawer and put the bowl inside.
5. Put the bowl on top of the cabinet.
6. Push the plate to the front of the stove.
7. Put the cream cheese in the bowl.
8. Turn on the stove.
9. Put the bowl on the plate.
10. Put the wine bottle on the rack.
The 10 tasks in LIBERO-10 are:
1. LIVING ROOM SCENE2: Put both the alphabet soup and the tomato sauce in the basket.
2. LIVING ROOM SCENE2: Put both the cream cheese box and the butter in the basket.
3. KITCHEN SCENE3: Turn on the stove and put the moka pot on it.
4. KITCHEN SCENE4: Put the black bowl in the bottom drawer of the cabinet and close it.
5. LIVING ROOM SCENE5: Put the white mug on the left plate and put the yellow and white mug on the right plate.
6. STUDY SCENE1: Pick up the book and place it in the back compartment of the caddy.
7. LIVING ROOM SCENE6: Put the white mug on the plate and put the chocolate pudding to the right of the plate.
8. LIVING ROOM SCENE1: Put both the alphabet soup and the cream cheese box in the basket.
9. KITCHEN SCENE8: Put both moka pots on the stove.
10. KITCHEN SCENE6: Put the yellow and white mug in the microwave and close it.
The 10 tasks in LIBERO-90 are:
1. KITCHEN SCENE10: Close the top drawer of the cabinet.
2. KITCHEN SCENE10: Close the top drawer of the cabinet and put the black bowl on top of it.
3. KITCHEN SCENE10: Put the black bowl in the top drawer of the cabinet.
4. KITCHEN SCENE10: Put the butter at the back in the top drawer of the cabinet and close it.
5. KITCHEN SCENE10: Put the butter at the front in the top drawer of the cabinet and close it.
6. KITCHEN SCENE10: Put the chocolate pudding in the top drawer of the cabinet and close it.
7. KITCHEN SCENE1: Open the bottom drawer of the cabinet.
8. KITCHEN SCENE1: Open the top drawer of the cabinet.
9. KITCHEN SCENE1: Open the top drawer of the cabinet and put the bowl in it.
10. KITCHEN SCENE1: Put the black bowl on the plate.
11. KITCHEN SCENE1: Put the black bowl on top of the cabinet.
12. KITCHEN SCENE2: Open the top drawer of the cabinet.
13. KITCHEN SCENE2: Put the black bowl at the back on the plate.
14. KITCHEN SCENE2: Put the black bowl at the front on the plate.
15. KITCHEN SCENE2: Put the middle black bowl on the plate.
16. KITCHEN SCENE2: Put the middle black bowl on top of the cabinet.
17. KITCHEN SCENE2: Stack the black bowl at the front on the black bowl in the middle.
18. KITCHEN SCENE2: Stack the middle black bowl on the back black bowl.
19. KITCHEN SCENE3: Put the frying pan on the stove.
20. KITCHEN SCENE3: Put the moka pot on the stove.
21. KITCHEN SCENE3: Turn on the stove.
22. KITCHEN SCENE3: Turn on the stove and put the frying pan on it.
23. KITCHEN SCENE4: Close the bottom drawer of the cabinet.
24. KITCHEN SCENE4: Close the bottom drawer of the cabinet and open the top drawer.
25. KITCHEN SCENE4: Put the black bowl in the bottom drawer of the cabinet.
26. KITCHEN SCENE4: Put the black bowl on top of the cabinet.
27. KITCHEN SCENE4: Put the wine bottle in the bottom drawer of the cabinet.
28. KITCHEN SCENE4: Put the wine bottle on the wine rack.
29. KITCHEN SCENE5: Close the top drawer of the cabinet.
30. KITCHEN SCENE5: Put the black bowl in the top drawer of the cabinet.
31. KITCHEN SCENE5: Put the black bowl on the plate.
32. KITCHEN SCENE5: Put the black bowl on top of the cabinet.
33. KITCHEN SCENE5: Put the ketchup in the top drawer of the cabinet.
34. KITCHEN SCENE6: Close the microwave.
35. KITCHEN SCENE6: Put the yellow and white mug to the front of the white mug.
36. KITCHEN SCENE7: Open the microwave.
37. KITCHEN SCENE7: Put the white bowl on the plate.
38. KITCHEN SCENE7: Put the white bowl to the right of the plate.
39. KITCHEN SCENE8: Put the right moka pot on the stove.
40. KITCHEN SCENE8: Turn off the stove.
41. KITCHEN SCENE9: Put the frying pan on the cabinet shelf.
42. KITCHEN SCENE9: Put the frying pan on top of the cabinet.
43. KITCHEN SCENE9: Put the frying pan under the cabinet shelf.
44. KITCHEN SCENE9: Put the white bowl on top of the cabinet.
45. KITCHEN SCENE9: Turn on the stove.
46. KITCHEN SCENE9: Turn on the stove and put the frying pan on it.
47. LIVING ROOM SCENE1: Pick up the alphabet soup and put it in the basket.
48. LIVING ROOM SCENE1: Pick up the cream cheese box and put it in the basket.
49. LIVING ROOM SCENE1: Pick up the ketchup and put it in the basket.
50. LIVING ROOM SCENE1: Pick up the tomato sauce and put it in the basket.
51. LIVING ROOM SCENE2: Pick up the alphabet soup and put it in the basket.
52. LIVING ROOM SCENE2: Pick up the butter and put it in the basket.
53. LIVING ROOM SCENE2: Pick up the milk and put it in the basket.
54. LIVING ROOM SCENE2: Pick up the orange juice and put it in the basket.
55. LIVING ROOM SCENE2: Pick up the tomato sauce and put it in the basket.
56. LIVING ROOM SCENE3: Pick up the alphabet soup and put it in the tray.
57. LIVING ROOM SCENE3: Pick up the butter and put it in the tray.
58. LIVING ROOM SCENE3: Pick up the cream cheese and put it in the tray.
59. LIVING ROOM SCENE3: Pick up the ketchup and put it in the tray.
60. LIVING ROOM SCENE3: Pick up the tomato sauce and put it in the tray.
61. LIVING ROOM SCENE4: Pick up the black bowl on the left and put it in the tray.
62. LIVING ROOM SCENE4: Pick up the chocolate pudding and put it in the tray.
63. LIVING ROOM SCENE4: Pick up the salad dressing and put it in the tray.
64. LIVING ROOM SCENE4: Stack the left bowl on the right bowl and place them in the tray.
65. LIVING ROOM SCENE4: Stack the right bowl on the left bowl and place them in the tray.
66. LIVING ROOM SCENE5: Put the red mug on the left plate.
67. LIVING ROOM SCENE5: Put the red mug on the right plate.
68. LIVING ROOM SCENE5: Put the white mug on the left plate.
69. LIVING ROOM SCENE5: Put the yellow and white mug on the right plate.
70. LIVING ROOM SCENE6: Put the chocolate pudding to the left of the plate.
71. LIVING ROOM SCENE6: Put the chocolate pudding to the right of the plate.
72. LIVING ROOM SCENE6: Put the red mug on the plate.
73. LIVING ROOM SCENE6: Put the white mug on the plate.
74. STUDY SCENE1: Pick up the book and place it in the front compartment of the caddy.
75. STUDY SCENE1: Pick up the book and place it in the left compartment of the caddy.
76. STUDY SCENE1: Pick up the book and place it in the right compartment of the caddy.
77. STUDY SCENE1: Pick up the yellow and white mug and place it to the right of the caddy.
78. STUDY SCENE2: Pick up the book and place it in the back compartment of the caddy.
79. STUDY SCENE2: Pick up the book and place it in the front compartment of the caddy.
80. STUDY SCENE2: Pick up the book and place it in the left compartment of the caddy.
81. STUDY SCENE2: Pick up the book and place it in the right compartment of the caddy.
82. STUDY SCENE3: Pick up the book and place it in the front compartment of the caddy.
83. STUDY SCENE3: Pick up the book and place it in the left compartment of the caddy.
84. STUDY SCENE3: Pick up the book and place it in the right compartment of the caddy.
85. STUDY SCENE3: Pick up the red mug and place it to the right of the caddy.
86. STUDY SCENE3: Pick up the white mug and place it to the right of the caddy.
87. STUDY SCENE4: Pick up the book in the middle and place it on the cabinet shelf.
88. STUDY SCENE4: Pick up the book on the left and place it on top of the shelf.
89. STUDY SCENE4: Pick up the book on the right and place it on the cabinet shelf.
90. STUDY SCENE4: Pick up the book on the right and place it under the cabinet shelf.